OpenStudy (anonymous):

Eliminate the parameter. x = t - 3, y = t2 + 5 a. y = x2 + 6x + 14 b. y = x2 - 14 c. y = x2 - 6x - 14 d. y = x2 + 14

OpenStudy (raffle_snaffle):

I don't really understand the question.. It's all you loser66

OpenStudy (anonymous):

t=x-3??

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[x=t-3\] solve for \(t\)

OpenStudy (loser66):

@thinktallthoughts you have t = x +3 not x -3, right?

OpenStudy (anonymous):

oh yeah.

OpenStudy (anonymous):

so then it's y=(x+3)2+5??

OpenStudy (loser66):

now, just replace that t on y = t^2 +5 DAT SIT

OpenStudy (loser66):

Yup

OpenStudy (anonymous):

thx!

OpenStudy (loser66):

:)